OpenStudy (anonymous):

Provide the solution to the equation 25 = sqrt 5(2x + 6).

OpenStudy (anonymous):

\[25= \sqrt{5(2x+6)}\] this?

OpenStudy (anonymous):

^2x + 6

OpenStudy (anonymous):



OpenStudy (anonymous):

\[25= \sqrt{5}^{2x+6}\]

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

\[25= 5^2\] \[sqrt{5}=5^{\frac{1}{2}}\]

OpenStudy (anonymous):

That's all you do?

OpenStudy (anonymous):

so \[25= \sqrt{5}^{2x+6}\] becomes \[5^2=5^{\frac{2x+6}{2}}\] this means \[2=\frac{2x+6}{2}\] and solve for x @lopez_hatesmath please dont just give people the answer without either explaining it first or if its a simple one step problem

OpenStudy (anonymous):

i figured you had it handled anyways man but thanks for the comment :)